# Graph-Informed Adversarial Modeling: Infimal Subadditivity of Interpolative Divergences

### 20 Mar 2026 v1

@@ Abstract
We study adversarial learning when the target distribution factorizes according to a known Bayesian network. For interpolative divergences, including $(f,\Gamma)$-divergences, we prove a new infimal subadditivity principle showing that, under suitable conditions, a global variational discrepancy is controlled by an average of family-level discrepancies aligned with the graph. In an additive regime, this surrogate is exact. This provides a variational justification for replacing a graph-agnostic GAN with a monolithic discriminator by a graph-informed GAN with localized family-level discriminators. The result does not require the optimizer itself to factorize according to the graph. We also obtain parallel results for integral probability metrics and proximal optimal transport divergences, identify natural discriminator classes for which the theory applies, and present experiments showing improved stability and structural recovery relative to graph-agnostic baselines.
@@ 1 Introduction
GANs are often trained without explicitly exploiting known conditional dependence structure in the target distribution. In this paper, we refer to such approaches as graph-agnostic when the discriminator acts on the full joint distribution and does not use a known graphical model constructed from side information. In some applications, however, side information about the distribution is available, such as known physical/chemical laws, covariates, design constraints, and expert judgments. In Feizi et al. (2017) the authors proposed GAN architectures tailored to multivariate Gaussian distributions, while Balaji et al. (2019) and Farnia et al. (2023) extended the analysis to GANs designed for mixtures of Gaussians. Birrell et al. (2022b) introduces structure-preserving GANs which leverage additional structure, such as group symmetry, to learn data-efficiently via new variational representations of divergences. The approach reduces the discriminator space to its projection onto the invariant subspace using conditional expectation with respect to the $\sigma$-algebra defined by the structure.
However, in this work we shift our focus to cases where conditional dependencies of variables involved in the model could be represented by a Bayesian network (see Pearl (1988); Koller and Friedman (2009)), e.g., the variables might follow a Markov chain. Formally, a Bayesian network is a joint probability distribution on a collection of random variables indexed by the vertices of a Directed Acyclic Graph (DAG). Let $G=(V,E)$ be a DAG with vertex set $V=\{1,\dots,n\}$, $n\in\mathbb{N}$, and edge set $E\subseteq V\times V$. We associate to each vertex $i\in V$ a random variable $X_{i}$ taking values in a space $\mathcal{X}_{i}$, and write $\bm{X}=\bm{X}_{V}=(X_{1},\dots,X_{n})$ with state space $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$. The Bayesian network is then the pair $(G,P)$, where the joint distribution $P$ of $\bm{X}$ factorizes according to $G$ as
$P(dx_{1},\dots,dx_{n})=\prod_{i=1}^{n}P_{i|{\mathrm{Pa}}(i)}\bigl(dx_{i}\,\big|\,\bm{X}_{\mathrm{Pa}(i)}=\bm{x}_{\mathrm{Pa}(i)}\bigr)\,$,
where $\mathrm{Pa}(i)$ denotes the parents of vertex $i$ and $P_{i\mid\mathrm{Pa}(i)}$ are the corresponding conditional probability distributions (CPDs) or marginal distributions if $\mathrm{Pa}(i)=\emptyset$. Intuitively, a Bayesian network encodes which variables directly depend on which others, providing a structured factorization of the joint distribution. In this paper, we refer to GANs that explicitly use this graphical information in the design of the discriminator objective as graph-informed.
@@
Motivated by these ideas, we establish an analogous variant of the subadditivity property on Bayesian networks, referred to as $\alpha$-linear generalized infimal subadditivity. More precisely, for a measure of discrepancy $\delta$ with respect to measures of discrepancy $\delta_{1},\dots,\delta_{n}$
@@
for all Bayesian networks $Q$, where $Q_{i\,\cup\,\mathrm{Pa}(i)}$ $P_{i\,\cup\,\mathrm{Pa}(i)}$ are the marginals of $Q$ and $P$ over the $i$-th node and its parents. We show that $(f,\Gamma)$-divergences and other interpolative divergences satisfy $\alpha$-linear generalized infimal subadditivity, whereas classical $f$-divergences satisfy $\frac{1}{n}$-linear generalized infimal additivity, that is, the case where the above inequality holds with equality.
In generative modeling, a monolithic discriminator acting on a high-dimensional, joint distribution poses both computational and statistical challenges: min–max optimization is costly, and hypothesis testing in high dimensions requires sample complexity that grows exponentially with the dimension. Guided by our main theoretical results, which establish $\alpha$-linear generalized infimal (sub)additivity principles for classes of statistical discrepancies, we propose graph-informed training strategies. Our work shows that, when the target has a known Bayesian network structure, a graph-informed GAN can replace a graph-agnostic GAN, achieving computational and statistical advantages.
This paper is organized as follows. In the next section, we place our contribution in the broader literature on structure-aware generative modeling and adversarial learning. In Section 3, we recall the variational approach for learning adversarial generative models and key statistical discrepancies. We introduce localized function classes and conditional expectation operators used to encode the factorization of the target probability distribution with respect to a Bayesian network in Section 4. Our main theoretical results, namely Theorems 5.4, 5.6, and 5.7, which prove infimal subadditivity for $(f,\Gamma)$-divergences, IPMs, and optimal transport divergences, are presented in Section 5. These provide guarantees (certificates) for replacing global discrepancy minimization by a principled, graph-informed surrogate. Finally, we present several numerical case studies in Section 6 and discuss implications and future work in Section 7.
@@ 2 Related work
These prior studies focus primarily on incorporating structure through the generator, through auxiliary constraints, or through the overall training architecture. That is, they focus on injecting structure through model design. By contrast, our contribution is on the discriminator-side and variational: a general variational principle grounded in a new infimal subadditivity result for $(f,\Gamma)$-divergences on distributions with known Bayesian network structure. We study how known Bayesian network structure can be used to replace a graph-agnostic monolithic discriminator objective by graph-informed family-level objectives, and we show that this yields a principled surrogate for the corresponding graph-constrained global discrepancy. Methodologically, this connects to empirical multicritic strategies used, for example, in image-to-image translation methods (Yu et al., 2019), recurrent neural networks and time-series generation (Hyland et al., 2018), convolutional neural networks (Nie et al., 2018), and inpainting (Chang et al., 2019). However, our results supply the missing theoretical rationale for these largely heuristic approaches: when the target factorizes according to a Bayesian network, sums of local graph-informed discrepancies control the corresponding global objective.
@@
Our contribution is not to introduce another architecture-specific mechanism for injecting structure, but to develop a general variational principle for graph-informed adversarial learning. To make this precise, we next review the variational formulation of GAN training and the discrepancy functionals that underpin our later subadditivity and infimal subadditivity results.
@@ 3.2 $f$-Divergences
where $f^{*}(s)=\sup_{t\in\mathbb{R}}\left\{st-f(t)\right\}$ is the Legendre–Fenchel transform (convex conjugate) of $f$. For variational training of GANs, the choice of $f$ determines both the divergence $\mathfrak{D}_{f}(Q\|P)$ and the nonlinearity in the discriminator through its convex conjugate $f^{*}$. In particular, the growth of $f^{*}$ controls how strongly large values $\gamma(x)$ are penalized under the reference distribution $Q$.
In the sequel, we will use the KL divergence with $f_{\mathrm{KL}}(t)=t\log t-t+1$ and the JS divergence with
$f_{\mathrm{JS}}(t)=\tfrac{1}{2}\left[t\log t-(t+1)\log\left(\frac{t+1}{2}\right)\right]\,,\qquad t\geq 0\,.$
The corresponding Legendre–Fenchel transforms are $f^{*}_{\mathrm{KL}}(s)=e^{s-1}$ and $f^{*}_{\mathrm{JS}}(s)=-\log(2-e^{s})$, $s<\log 2$ (see, e.g., Nowozin et al. (2016)). In the KL case, $f^{*}_{\mathrm{KL}}$ has exponential growth, which heavily penalizes large discriminator values and emphasizes rare events where $P$ and $Q$ disagree in the tails. Moreover, for the KL divergence the infimum over $\nu$ in equation 4 can be solved analytically and one recovers the classical Donsker–Varadhan representation (cf. equation 34), in which the discriminator objective involves a log-moment generating function $\log E_{P}[e^{\gamma}]$ term, a form that underlies many existing $f$-GAN and risk-sensitive objectives. Unlike KL, the JS divergence is symmetric and bounded, $0\leq\mathfrak{D}_{\mathrm{JS}}(Q\|P)\leq\log 2$, and the corresponding conjugate $f^{*}_{\mathrm{JS}}$ diverges only logarithmically as $s\to\log 2$, so the penalty on extreme discriminator values is much milder than when using KL. In the JS case, the variational representation recovers the original logistic GAN objective (up to an additive constant and a reparameterization of $\gamma$). In our experiments involving discrete Bayesian networks, we therefore focus on the JS divergence, since its symmetry and boundedness (it remains finite even when $P$ and $Q$ have partially mismatched support) tend to yield more stable training than KL-based objectives in this setting.
@@ 3.3 $\Gamma$-Integral Probability Metrics ($\Gamma$-IPMs)
We next turn to $\Gamma$-IPMs, which arise when the discriminator is constrained to a prescribed function class and include several widely used probability metrics. Let $\Gamma\subset\mathcal{M}_{b}(\mathcal{X})$, a class of bounded measurable test functions on $\mathcal{X}$, and the associated $\Gamma$-IPM as,
@@ 3.5 Proximal optimal transport divergences
Finally, we consider proximal optimal transport divergences, which provide an interpolation between transport-based and information-theoretic discrepancies and fit naturally into the variational framework developed above. Proximal optimal transport divergence, introduced in Baptista et al. (2025), is a discrepancy measure that interpolates between information divergences and optimal transport distances via an infimal convolution formulation,
$\mathfrak{D}^{c}_{\varepsilon}(Q\|P)=\inf_{R}\left\{T_{c}(Q,R)+\varepsilon\mathfrak{D}_{f}(R\|P)\right\}\,$, (9)
where the infimum is over all probability distributions $R$ on a Polish space $\mathcal{X}$ and $T_{c}$ is the optimal transport cost corresponding to a cost function $c:\mathcal{X}\times\mathcal{X}\to[0,\infty)$. When $T_{c}$ is the $1$-Wasserstein distance, then equation 9 becomes an $(f,\Gamma)$-divergence with $f=f_{{\rm{KL}}}$ and $\Gamma={\rm{Lip}}^{1}_{b}(\mathcal{X})$. As in $(f,\Gamma)$-divergence the dual formulation offers a variational representation involving test functions. For example, for $f=f_{{\rm{KL}}}$ and $c$ bounded below and lower semicontinuous, the dual representation is given by
$\mathfrak{D}_{\mathrm{KL},\varepsilon}^{c}(Q\|P)=\sup_{\begin{subarray}{c}(\phi,\psi)\in C_{b}(\mathcal{X})\times C_{b}(\mathcal{X}),\\ \phi\oplus\psi\leq c\end{subarray}}\left\{E_{Q}[\phi]-\varepsilon\ln E_{P}[e^{-\psi/\varepsilon}]\right\},$ (10)
where $\phi\oplus\psi$ denotes the function $\phi(x)+\psi(y)$. For general cost functions $c$, we cannot expect the optimizers of the optimal transport cost to be bounded, but there exist optimizers $\phi\in L^{1}(Q)$ and $\psi\in L^{1}(P)$ such that $\phi\oplus\psi\leq c$.
@@ 4 Bayesian network operators
Continuing the Bayesian network notation from Section 1, for a DAG denoted $G=(V,E)$, the random vector $\bm{X}=\bm{X}_{V}=(X_{1},\dots,X_{n})$ indexed by the vertices $V$ takes values $\bm{x}=(x_{1},\dots,x_{n})$ in the state space $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$. We denote the joint probability distribution of $\bm{X}$ by $Q$ (with density $q$ if it exists). For any subset $A=\{i_{1},\dots,i_{m}\}\subset V$, we denote $\bm{X}_{A}=(X_{i_{1}},\dots,X_{i_{m}})$ which takes values $\bm{X}_{A}=\bm{x}_{A}=(x_{i_{1}},\dots,x_{i_{m}})\in\mathcal{X}_{A}=\mathcal{X}_{i_{1}}\times\cdots\times X_{i_{m}}$ and its marginal is denoted by $Q_{A}$. The conditional probability distribution of $X_{i}$ given parents values $\bm{X}_{\mathrm{Pa}(i)}=\bm{x}_{\mathrm{Pa}(i)}$ is denoted by $Q_{i\mid\mathrm{Pa}(i)}$, i.e, $Q_{i\mid\mathrm{Pa}(i)}(dx_{i})=Q(dx_{i}\mid\bm{X}_{\mathrm{Pa}(i)}=\bm{x}_{\mathrm{Pa}(i)})$ with corresponding density $q(x_{i}\mid\bm{x}_{\mathrm{Pa}(i)})$. In general, for two sets of indices $A,B\subset V$, $Q_{A\mid B}(d\bm{x}_{A})=Q(d\bm{x}_{A}\mid\bm{X}_{B}=\bm{x}_{B})$. Having all notation and terminology in place, we next define Bayesian networks formally.
@@
$Q(dx_{1},\dots,dx_{n})=\prod_{i=1}^{n}Q_{i\mid\mathrm{Pa}(i)}(dx_{i}\mid\bm{X}_{\mathrm{Pa}(i)}=\bm{x}_{\mathrm{Pa}(i)})\,$, (11)
@@
We now fix $Q\in\mathcal{P}^{G}$. For each vertex $i\in V$, we define the following spaces of probability distributions:
@@
$I_{i,{\rm{Pa}}(i)}[\gamma](x_{i},{\bm{x}}_{{\rm{Pa}}_{\,i}})\!=\!\!\int_{{\bm{\mathcal{X}}}_{V^{(i-1)}}}\!\!\int_{\mathcal{X}_{i+1}}\!\!\!\!\!\cdots\!\int_{\mathcal{X}_{n}}\!\!\!\!\gamma({\bm{x}})\!\!\prod_{j=i+1}^{n}\!\!Q_{j\mid\mathrm{Pa}(j)}(dx_{j})\,Q_{V^{(i-1)}\mid\mathrm{Pa}(i)}(d{\bm{x}}_{V^{(i-1)}})\,.$ (13)
@@
5 Graph-informed GANs
In this section, we combine the variational discrepancy framework of Section 3 with the Bayesian-network operators introduced in Section 4. We show that the global discrepancy minimization can be replaced by a principled surrogate, a minimization over family-level objectives on the neighborhoods $i\cup\mathrm{Pa}(i)$. We first recall generalized subadditivity from Ding et al. (2021), then introduce an infimal version suited to interpolative divergences, and finally state our main results for $(f,\Gamma)$-divergences, IPMs, and proximal optimal transport divergences. These results provide the theoretical basis for graph-informed GANs.
In Ding et al. (2021), the authors establish generalized subadditivity bounds for a broad class of probability distances and divergences, including the Jensen–Shannon divergence, the Wasserstein distance, and nearly all $f$-divergences. We reproduce these key bounds in Definition 5.1 below for the convenience of the reader. Throughout the remainder of this section, we fix a graph structure $G$ and recall that $\mathcal{P}^{G}$ denotes the class of probability distributions that factorize according to $G$.

### 02 Apr 2026 v2

@@ Abstract
We study adversarial learning when the target distribution factorizes according to a known Bayesian network. For interpolative divergences, including $(f,\Gamma)$-divergences, we prove a new infimal subadditivity principle showing that, under suitable conditions, a global variational discrepancy is controlled by an average of family-level discrepancies aligned with the graph. In an additive regime, the surrogate is exact. This closes a theoretical gap in the literature; existing subadditivity results justify graph-informed adversarial learning for classical discrepancies, but not for interpolative divergences, where the usual factorization argument breaks down. In turn, we provide a justification for replacing a standard, graph-agnostic GAN with a monolithic discriminator by a graph-informed GAN (GiGAN) with localized family-level discriminators, without requiring the optimizer itself to factorize according to the graph. We also obtain parallel results for integral probability metrics and proximal optimal transport divergences, identify natural discriminator classes for which the theory applies, and present experiments showing improved stability and structural recovery relative to graph-agnostic baselines.
@@ 1 Introduction
GANs are often trained without explicitly exploiting known conditional dependence structure in the target distribution. In this paper, we refer to such approaches as graph-agnostic when the discriminator acts on the full joint distribution and does not use a known graphical model constructed from side information. In some applications, however, side information about the distribution is available, such as known physical/chemical laws, covariates, design constraints, and expert judgments. This perspective fits within a broader line of work in which probabilistic graphical structure is used to inform neural network learning architectures and surrogate models, see, e.g. Hall et al. (2021). In Feizi et al. (2017) the authors proposed GAN architectures tailored to multivariate Gaussian distributions, while Balaji et al. (2019) and Farnia et al. (2023) extended the analysis to GANs designed for mixtures of Gaussians. Birrell et al. (2022b) introduces structure-preserving GANs which leverage additional structure, such as group symmetry, to learn data-efficiently via new variational representations of divergences. The approach reduces the discriminator space to its projection onto the invariant subspace using conditional expectation with respect to the $\sigma$-algebra defined by the structure.
However, in this work we shift our focus to cases where conditional dependencies of variables involved in the model could be represented by a Bayesian network (see Pearl (1988); Koller and Friedman (2009)), e.g., the variables might follow a Markov chain. Formally, a Bayesian network is a joint probability distribution on a collection of random variables indexed by the vertices of a Directed Acyclic Graph (DAG). Let $G=(V,E)$ be a DAG with vertex set $V=\{1,\dots,n\}$, $n\in\mathbb{N}$, and edge set $E\subseteq V\times V$. We associate to each vertex $i\in V$ a random variable $X_{i}$ taking values in a space $\mathcal{X}_{i}$, and write $\bm{X}=\bm{X}_{V}=(X_{1},\dots,X_{n})$ with state space $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$. The Bayesian network is then the pair $(G,Q)$, where the joint distribution $Q$ of $\bm{X}$ factorizes according to $G$ as
$Q(dx_{1},\dots,dx_{n})=\prod_{i=1}^{n}Q_{i|{\mathrm{Pa}}(i)}(dx_{i})\,$,
where $\mathrm{Pa}(i)$ denotes the parents of vertex $i$ and $Q_{i|\mathrm{Pa}(i)}=Q(dx_{i}|\bm{x}_{\mathrm{Pa}(i)})$ are the corresponding conditional probability distributions (CPDs) or marginal distributions if $\mathrm{Pa}(i)=\emptyset$. Intuitively, a Bayesian network encodes which variables directly depend on which others, providing a structured factorization of the joint distribution. In this paper, we refer to GANs that explicitly use this graphical information in the design of the discriminator objective as graph-informed.
@@
In generative modeling, a monolithic discriminator acting on a high-dimensional, joint distribution poses both computational and statistical challenges: the associated min–max optimization is costly, and discrimination via hypothesis testing in high dimensions can become increasingly sample-inefficient. Motivated by these issues, we establish an analogous variant of the subadditivity property on Bayesian networks, referred to as $\alpha$-linear generalized infimal subadditivity. More precisely, for a measure of discrepancy $\delta$ with respect to measures of discrepancy $\delta_{1},\dots,\delta_{n}$, this new property takes the form,
@@
for all Bayesian networks $Q$, where $Q_{i\,\cup\,\mathrm{Pa}(i)}$, $P_{i\,\cup\,\mathrm{Pa}(i)}$ are the marginals of $Q$ and $P$ over the $i$-th node and its parents. We show that $(f,\Gamma)$-divergences and other interpolative divergences satisfy $\alpha$-linear generalized infimal subadditivity, whereas classical $f$-divergences satisfy $\frac{1}{n}$-linear generalized infimal additivity, that is, the case where the above inequality holds with equality.
This extends the scope of graph-informed adversarial learning that can be justified. Existing results explain localized discriminators mainly through classical subadditivity properties of unrestricted discrepancies; by introducing infimal subadditivity, we show that an analogous justification continues to hold for interpolative divergences, including $(f,\Gamma)$-divergences, where the classical argument breaks down. In generative modeling, this means that when the target has a known Bayesian-network structure, a graph-informed GAN trained under an interpolative divergence objective (i.e., GiGAN) can replace a graph-agnostic GAN with a monolithic discriminator through a principled surrogate.
This paper is organized as follows. In the next section, we place our contribution in the broader literature on structure-aware generative modeling and adversarial learning. In Section 3, we recall the variational approach for learning adversarial generative models and key statistical discrepancies. We introduce localized function classes and conditional expectation operators used to encode the factorization of the target probability distribution with respect to a Bayesian network in Section 4. Our main theoretical results, namely Theorems 5.3, 5.5, and 5.6, which prove infimal subadditivity for $(f,\Gamma)$-divergences, IPMs, and optimal transport divergences, are presented in Section 5. These provide guarantees (certificates) for replacing global discrepancy minimization by a principled, graph-informed surrogate. Finally, we present several numerical case studies in Section 6 and discuss implications and future work in Section 7.
@@ 2 Related work
These prior studies focus primarily on incorporating structure through the generator, through auxiliary constraints, or through the overall training architecture. That is, they focus on injecting structure through model design. By contrast, our contribution is on the discriminator-side and variational. We study how known Bayesian network structure can be used to replace a graph-agnostic monolithic discriminator objective by graph-informed family-level objectives, and we show that this yields a principled surrogate for the corresponding graph-constrained global discrepancy. Methodologically, this connects to empirical multicritic strategies used, for example, in image-to-image translation methods (Yu et al., 2019), recurrent neural networks and time-series generation (Hyland et al., 2018), convolutional neural networks (Nie et al., 2018), and inpainting (Chang et al., 2019). However, our results supply the missing theoretical rationale for these largely heuristic approaches: when the target factorizes according to a Bayesian network, sums of local graph-informed discrepancies control the corresponding global objective.
@@
Our contribution is not to introduce another architecture-specific mechanism for injecting structure, but to develop a general variational principle for graph-informed adversarial learning. In contrast to prior work, which relied on classical subadditivity properties of unrestricted divergences, we show that interpolative divergences admit a weaker but still useful infimal subadditivity principle. This fills a theoretical gap in the literature and provides a general justification for graph-informed localized discriminators in variational generative modeling. To make this precise, we next review the variational formulation of GAN training and the discrepancy functionals that underpin our later subadditivity and infimal subadditivity results.
@@ 3.2 $f$-Divergences
where $f^{*}(s)=\sup_{t\in[0,\infty)}\left\{st-f(t)\right\}$ is the Legendre–Fenchel transform (convex conjugate) of $f$. For variational training of GANs, the choice of $f$ determines both the divergence $\mathfrak{D}_{f}(Q\|P)$ and the nonlinearity in the discriminator through its convex conjugate $f^{*}$. In particular, the growth of $f^{*}$ controls how strongly large values $\gamma(x)$ are penalized under the reference distribution $P$.
In the sequel, we will use the Donsker–Varadhan variational formula for KL,
$\mathfrak{D}_{\rm DV}(Q\|P)=\sup_{\gamma\in\mathcal{M}_{b}(\mathcal{X})}\left\{E_{Q}[\gamma]-\log E_{P}[e^{\gamma}]\right\}\,$, (5)
and, when considering $f$-GAN type losses, we will also use equation 4 with both $f_{\mathrm{KL}}(t)=t\log t$, for the KL divergence, and
$f_{\mathrm{JS}}(t)=\tfrac{1}{2}\left[t\log t-(t+1)\log\left(\frac{t+1}{2}\right)\right]\,,\qquad t\geq 0\,$
for the JS divergence. The corresponding Legendre–Fenchel transforms are $f^{*}_{\mathrm{KL}}(s)=e^{s-1}$ and $f^{*}_{\mathrm{JS}}(s)=-\log(2-e^{s})$, $s<\log 2$ (see, e.g., (Nowozin et al., 2016, Tables 1 and 2)). In the KL case, $f^{*}_{\mathrm{KL}}$ has exponential growth, which heavily penalizes large discriminator values and emphasizes rare events where $P$ and $Q$ disagree in the tails. Moreover, for the KL divergence the infimum over $\nu$ in equation 4 can be solved analytically and one recovers the classical Donsker–Varadhan representation equation 5, in which the discriminator objective involves a log-moment generating function, a form that underlies many existing $f$-GAN and risk-sensitive objectives. Unlike KL, the JS divergence is symmetric and bounded, $0\leq\mathfrak{D}_{\mathrm{JS}}(Q\|P)\leq\log 2$, and the corresponding conjugate $f^{*}_{\mathrm{JS}}$ diverges only logarithmically as $s\to\log 2$, so the penalty on extreme discriminator values is much milder than when using KL. In the JS case, the variational representation recovers the original logistic GAN objective (up to an additive constant and a reparameterization of $\gamma$). In our experiments involving discrete Bayesian networks, we therefore focus on the JS divergence, since its symmetry and boundedness (it remains finite even when $P$ and $Q$ have partially mismatched support) tend to yield more stable training than KL-based objectives in this setting.
@@ 3.3 $\Gamma$-Integral Probability Metrics ($\Gamma$-IPMs)
We next turn to $\Gamma$-IPMs, which arise when the discriminator is constrained to a prescribed function class and include several widely used probability metrics. Let $\Gamma\subset\mathcal{M}_{b}(\mathcal{X})$ be a class of bounded measurable test functions on $\mathcal{X}$, and define the associated $\Gamma$-IPM by,
@@ 3.5 Proximal optimal transport divergences
Finally, we consider proximal optimal transport divergences, which provide an interpolation between transport-based and information-theoretic discrepancies and fit naturally into the variational framework developed above. Proximal optimal transport divergence, introduced in Baptista et al. (2025), is a discrepancy measure that interpolates between information divergences $\mathfrak{D}$ and optimal transport distances $T_{c}$ corresponding to a cost function $c:\mathcal{X}\times\mathcal{X}\to[0,\infty)$, via an infimal convolution formulation,
$\mathfrak{D}^{c}_{\varepsilon}(Q\|P)=\inf_{R}\left\{T_{c}(Q,R)+\varepsilon\mathfrak{D}(R\|P)\right\}\,$, (10)
where the infimum is over all probability distributions $R$ on a Polish space $\mathcal{X}$. As for the $(f,\Gamma)$-divergence, the dual formulation offers a variational representation involving test functions. For example, for $f=f_{{\rm{KL}}}$ and $c$ bounded below and lower semicontinuous, the dual representation is given by
$\mathfrak{D}_{\mathrm{KL},\varepsilon}^{c}(Q\|P)=\sup_{\begin{subarray}{c}(\phi,\psi)\in C_{b}(\mathcal{X})\times C_{b}(\mathcal{X}),\\ \phi\oplus\psi\leq c\end{subarray}}\left\{E_{Q}[\phi]-\varepsilon\log E_{P}[e^{-\psi/\varepsilon}]\right\},$ (11)
where $\phi\oplus\psi$ denotes the function $\phi(x)+\psi(y)$. While the dual problem is written over bounded continuous functions for convenience, optimal potentials need not be bounded in general; however, there exist $\phi\in L^{1}(Q)$ and $\psi\in L^{1}(P)$ satisfying $\phi\oplus\psi\leq c$.
A special class of optimal transport distances on a metric space $(\mathcal{X},d)$ is when $T_{c}$ are $p$-Wasserstein distances,
$W_{p}(Q,P)=\inf_{\pi\in\Pi(Q,P)}\left(\int_{\mathcal{X}\times\mathcal{X}}d(x,y)^{p}\,d\pi(x,y)\right)^{\frac{1}{p}},$ (12)
where $c(x,y)=d(x,y)^{p}$ for $p\geq 1$, $Q,P\in\mathcal{P}_{p}(\mathcal{X})=\left\{P\in\mathcal{P}(\mathcal{X})\,:\,\int d(x,x_{0})^{p}dP<\infty\right\}$, and $x_{0}$ is an arbitrary point in $\mathcal{X}$. When $T_{c}$ is the $1$-Wasserstein distance, then equation 10 becomes an $(f,\Gamma)$-divergence with $f=f_{{\rm{KL}}}$ and $\Gamma={\rm{Lip}}^{1}_{b}(\mathcal{X})$.
@@ 4 Bayesian network operators
Continuing the Bayesian network notation from Section 1, for a DAG denoted $G=(V,E)$, the random vector $\bm{X}=\bm{X}_{V}=(X_{1},\dots,X_{n})$ indexed by the vertices $V$ takes values $\bm{x}=(x_{1},\dots,x_{n})$ in the state space $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$. We denote the joint probability distribution of $\bm{X}$ by $Q$ (with density $q$ if it exists). For any subset $A=\{i_{1},\dots,i_{m}\}\subset V$, we denote $\bm{X}_{A}=(X_{i_{1}},\dots,X_{i_{m}})$ which takes values $\bm{X}_{A}=\bm{x}_{A}=(x_{i_{1}},\dots,x_{i_{m}})\in\mathcal{X}_{A}=\mathcal{X}_{i_{1}}\times\cdots\times X_{i_{m}}$ and its marginal is denoted by $Q_{A}$. The conditional probability distribution of $X_{i}$ given parents values $\bm{X}_{\mathrm{Pa}(i)}=\bm{x}_{\mathrm{Pa}(i)}$ is denoted by $Q_{i\mid\mathrm{Pa}(i)}$, i.e, $Q_{i\mid\mathrm{Pa}(i)}(dx_{i})=Q(dx_{i}\mid\bm{x}_{\mathrm{Pa}(i)})$ with corresponding density $q(x_{i}|\bm{x}_{\mathrm{Pa}(i)})$. In general, for two sets of indices $A,B\subset V$, $Q_{A\mid B}(d\bm{x}_{A})=Q(d\bm{x}_{A}|\bm{x}_{B})$. For notational simplicity, when the index sets are clear, we write $Q_{i_{1}\cdots i_{m}}$ in place of $Q_{\{i_{1},\dots,i_{m}\}}$ and $Q_{i_{1}\cdots i_{m}\mid j_{1}\cdots j_{k}}$ in place of $Q_{\{i_{1},\dots,i_{m}\}\mid\{j_{1},\dots,j_{k}\}}$, e.g., in the proof of the main theorem. Having all notation and terminology in place, we next define Bayesian networks formally.
@@
$Q(dx_{1},\dots,dx_{n})=\prod_{i=1}^{n}Q_{i\mid\mathrm{Pa}(i)}(dx_{i})\,$, (13)
@@
We now fix $Q\in\mathcal{P}^{G}$. For each vertex $i\in V$, we the following spaces of probability distributions:
@@
$I_{i,{\rm{Pa}}(i)}[\gamma](x_{i},{\bm{x}}_{\mathrm{Pa}(i)})\!=\!\!\int_{{\bm{\mathcal{X}}}_{V^{(i-1)}}}\!\!\int_{\mathcal{X}_{i+1}}\!\!\!\!\!\cdots\!\int_{\mathcal{X}_{n}}\!\!\!\!\gamma({\bm{x}})\!\!\prod_{j=i+1}^{n}\!\!Q_{j\mid\mathrm{Pa}(j)}(dx_{j})\,Q_{V^{(i-1)}\mid\mathrm{Pa}(i)}(d{\bm{x}}_{V^{(i-1)}})\,.$ (15)
@@
5 Graph-informed GANs (GiGANs)
In this section, we combine the variational discrepancy framework of Section 3 with the Bayesian-network operators introduced in Section 4. Our goal is to show that, when the target distribution factorizes according to a known DAG, global discrepancy minimization can be replaced by a principled surrogate involving family-level objectives on the local neighborhoods $i\cup\mathrm{Pa}(i)$. We begin by recalling the generalized subadditivity framework of Ding et al. (2021), then explain why this classical notion is not well suited to interpolative divergences, and finally introduce the weaker infimal formulation that underpins our main results for $(f,\Gamma)$-divergences, IPMs, and proximal optimal transport divergences.
In Ding et al. (2021), the authors establish generalized subadditivity bounds for a broad class of probability distances and divergences, including the Jensen–Shannon divergence, Wasserstein distance, and nearly all $f$-divergences. For convenience, we recall their definition below. Throughout this section, we fix a graph structure $G$ write $\mathcal{P}^{G}$ for the class of probability distributions that factorize according to $G$.